HUBBLE SPACE TELESCOPE DAILY REPORT #1057 PERIOD COVERED: 0700 EST Daily Status Report as of 020/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed S/C 4721 (WF/PC-2 Decontamination and TEC Turn On) The WF/PC-2 was used to make ultraviolet photometric observations of the star standard A+81D262, plus K-Spots, and dark calibration images. All four sets of CCDs were used. A total of 11 images were received and displayed in the OSS. One of the images was reported to have an unexpected bias level change in a single chip. The other images were reported to be okay. Also, there was an East Coast power emergency outage which effected the science cluster computers. They were successfully brought back up later. 1.2 Completed S/C 5254 (GHRS SMOV Side One Activation) The Side One DEB was successfully turned on, at three hours intervals. The HRS was safed and recovered as scheduled with on problems noted. 1.3 Completed COSTAR/U 4747 (COSTAR/FOS Fine Alignment - X Tilt/Y tilt) Both the Faint Object Spectrograph and the COSTAR were used to perform alignment and focus operations on the FOS M1 mirror. The observations were received and displayed in the OSS. They were reported to look okay. The uplinks to change the focus have been canceled because the required preceding data has not been obtained. 1.4 Completed FGS 4892 (Duplicity Among Hyades Stars, Cycle 3 High) The FGSs were used to make astrometric observations of the possible binary star VA334 which is located in the Hyades open cluster. The observations were made in fine lock. The OSS monitored the observation, and the astrometric observations were reported to look fine. 1.5 Completed WF/PC-2 5240 (SMOV-WF/PC-2 Flat Field Calibration) The WF/PC-2 was used to make a set of visual internal flat images. Six images were received and displayed. One image was reported to be underexposed, and the others were okay. 1.6 Completed WF/PC-2 4709 (WF/PC-2 Internal Monitor) The Faint Object Spectrograph (Blue and Red Sides) was used to make ten spectral observations of cluster NGC-199 to help determine the location of the 4.3 arcsecond aperture. The observations were received and displayed. Two observations were underexposed, two were saturated, and the others were reported to be okay. 1.7 Completed Two Sets of WF/PC-2 4745 (COSTAR/FOS Fine Alignment - Focus) Both the Faint Object spectrograph and the COSTAR were were scheduled as part of a alignment and focus procedure of the FOS M1 mirror. The uplinks to change the focus were canceled because the required preceding data had not yet been obtained. 1.8 Completed GHRS/2 5549 (GHRS Side 2 Dark Noise Monitor - Cycle 4) The High Resolution spectrograph was used to make routine test to monitor dark noise in the detector. The observations were monitored in the OSS, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Goddard Space Flight Center was closed today because of the weather. No report was submitted. 2.2 FHST Updates: Goddard Space Flight Center was closed today because of the weather. No report was submitted. 2.3 Significant HSTARs: None 2.4 Operations Notes: There was an East Coast power emergency outage which was reported to have effected the science cluster computers at the STScI at several different times during the evening. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOS/BL 4745 (COSTAR/FOS Fine Alignment - Focus) 3.2 Complete S/C 5254 (GHRS SMOV Side 1 Activation) 3.3 Complete WF/PC-2 4763 (WF/PC-2 QE and Throughput Measurement) 3.4 Complete Two Sets of WF/PC-2 5240 (SMOV-WF/PC-2 Flat Fields Calibration) 3.5 Complete FOC/96 4788 (SMOV COSTAR/FOC Point Spread Function) 3.6 Complete COSTAR/U 4814 (COSTAR GHRS M2 Mirror Arm Deployment) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: None /DMH -END-